OpenStudy (anonymous):

the sum of the measure of the interior angle of a triangle is equal to_____. A. 90 B.180 C.270 D.360

OpenStudy (jadeishere):

So all 3 interior angles in a triangle add up to 180 >.< I don't actually know why, though >.< @Elsa213

OpenStudy (michele_laino):

hint: the sum S of internal angles of a polygon is given by the subsequent formula: \[\Large S = \left( {n - 2} \right) \times 180\] where n is the number of sides of our polygon. In our case we have a triangle, so, n=3

OpenStudy (michele_laino):

then, after a substitution, we get: \[\Large S = \left( {n - 2} \right) \times 180 = \left( {3 - 2} \right) \times 180 = ...?\]

OpenStudy (anonymous):

90 is the answer